A disk of radius R has positive charge uniformly distributed over a inner circular region of radius A and negative charge uniformly distributed over the outer annular region. The surface density on the inner region is +o and that of the outer is - o. The electrostatic potential at P located on the central perpendicular axis is a distance z=R from the disk center is zero

What is A? In term of R
1
Expert's answer
2021-10-18T11:01:19-0400

The potential from the outer ring:

"V_p=\frac\sigma{2\epsilon_0}[R\sqrt{2}-\sqrt{R^2+(R-A)^2}]."

The potential of the inner disk:


"V_p=\frac\sigma{2\epsilon_0}(R\sqrt{2}-R)."


For the field to be zero, these potentials must be equal:


"\frac\sigma{2\epsilon_0}[R\sqrt{2}-\sqrt{R^2+(R-A)^2}]=\frac\sigma{2\epsilon_0}[R\sqrt{2}-R],\\\space\\
\sqrt{R^2+(R-A)^2}=R,\\
R^2+R^2-A^2-2AR=R^2,\\
A^2+2AR-R^2=0,\\\space\\
A=\frac{\sqrt{4R^2+4R^2}-2R}{2}=R(\sqrt2-1)."